Alpha and Beta are divisions within the same company. The managers of both divisions are evaluated based on their own division’s return on investment (ROI). Assume the following information relative to the two divisions:

Case
1 2 3 4
Alpha Division:
Capacity in units 55,000 309,000 108,000 210,000
Number of units now being sold to
outside customers
55,000 309,000 83,000 210,000
Selling price per unit to outside
customers
$ 100 $ 40 $ 65 $ 43
Variable costs per unit $ 63 $ 20 $ 39 $ 27
Fixed costs per unit (based on
capacity)
$ 27 $ 9 $ 24 $ 4
Beta Division:
Number of units needed annually 10,400 69,000 20,000 58,000
Purchase price now being paid to
an outside supplier
$ 93 $ 39 $ 65 *

*Before any purchase discount.

Managers are free to decide if they will participate in any internal transfers. All transfer prices are negotiated.

Required:

1. Refer to case 1 shown above. Alpha Division can avoid $5 per unit in commissions on any sales to Beta Division.

a. What is the lowest acceptable transfer price from the perspective of the Alpha Division?

b. What is the highest acceptable transfer price from the perspective of the Beta Division?

c. What is the range of acceptable transfer prices (if any) between the two divisions? Will the managers probably agree to a transfer?

2. Refer to case 2 shown above. A study indicates that Alpha Division can avoid $3 per unit in shipping costs on any sales to Beta Division.

a. What is the lowest acceptable transfer price from the perspective of the Alpha Division?

b. What is the highest acceptable transfer price from the perspective of the Beta Division?

c. What is the range of acceptable transfer prices (if any) between the two divisions? Would you expect any disagreement between the two divisional managers over what the exact transfer price should be?

d. Assume Alpha Division offers to sell 69,000 units to Beta Division for $38 per unit and that Beta Division refuses this price. What will be the loss in potential profits for the company as a whole?

3. Refer to case 3 shown above. Assume that Beta Division is now receiving an 6% price discount from the outside supplier.

a. What is the lowest acceptable transfer price from the perspective of the Alpha Division?

b. What is the highest acceptable transfer price from the perspective of the Beta Division?

c. What is the range of acceptable transfer prices (if any) between the two divisions? Will the managers probably agree to a transfer?

d. Assume Beta Division offers to purchase 20,000 units from Alpha Division at $56.10 per unit. If Alpha Division accepts this price, would you expect its ROI to increase, decrease, or remain unchanged?

4. Refer to case 4 shown above. Assume that Beta Division wants Alpha Division to provide it with 58,000 units of a different product from the one Alpha Division is producing now. The new product would require $23 per unit in variable costs and would require that Alpha Division cut back production of its present product by 29,000 units annually. What is the lowest acceptable transfer price from Alpha Division’s perspective?
Answer #1

1.a.The Alpha division is operating at full capacity, the lowest acceptable transfer price from the perspective of the Alpha Division = Market Price – Costs Avoided

= $100-$5 = $95

b. highest acceptable transfer price from the perspective of the Beta Division = purchase price being paid to outside supplier

= $93

c.No range of acceptable transfer prices. No, managers will not agree for a transfer

2.a. The Alpha division is operating at full capacity, the lowest acceptable transfer price from the perspective of the Alpha Division = Market Price – Costs Avoided

= $40-$3 = $37          

b. highest acceptable transfer price from the perspective of the Beta Division = purchase price being paid to outside supplier

= $39

c.Range of Acceptable transfer price = $37-$39. Yes, managers will agree for a transfer> yes, there will be disagreement as both the managers will try to maximise their own profits

d.Loss in profits = 69000*(39-37) = $138,000

3.Spare capacity for alpha division

a.the lowest acceptable transfer price from the perspective of the Alpha Division = variable cost per unit

= $39

b. highest acceptable transfer price from the perspective of the Beta Division = purchase price being paid to outside supplier = 65*(1-6%) = $61.1

c.Range = $39-$61.1. Yes, ,managers will agree since benefit for both

d.ROI will increase since the transfer price is greater than variable cost and space capacity exists. Hence, transfer will lead to additional profits.

4.Lowest acceptable transfer price = cost incurred+benefit lost

= 23 + (43-27)/2

= $31 per unit